Fe2O3 + 3C = 2Fe + 3CO, 25.0 Mol Fe2O3, how many iron can be produced



1
Expert's answer
2021-04-28T06:17:38-0400

Fe2O3 can produce n mol of Fe;n =50 mol


Now you can calculate mass of Fe ( Mw of Fe =55.85)


"n=\frac {m} {Mw}"


m=n×Mw


m=50×55.85=2794 g